# Automatic Pruning via Structured Lasso with Class-wise Information

### v1

@@ 3.2.3 Tree-Guided Lasso and sTLP-IB
Where each group of $j$-th subtree regression factors $\beta_{j}^{G_{v}}$ is weighted by $w_{ve}$, as defined in (7), where $A$ is the ancestor set. $\beta_{j}^{G_{v}}$ is a vector and represents regression factors $\{\beta_{j}^{k}\mid k\in G_{v}\}$. Each node $v\in V$ of one tree is associated with a group $G_{v}$, whose members are the columns of $\beta$ within the nodes of the same subtree.
@@
Generally, $h_{v}$ represents the weight for selecting covariates relevant specifically to the column of $\beta$ linked to each child of node $v$, while $1-h_{v}$ is the weight for jointly selecting covariates relevant to all children of node $v$, $h_{v}\in(0,1)$. Assuming $|V|$ represents the number of nodes in a tree, since a tree associated with $out_{i}$ columns of $\beta$ for $C_{i}$ has $12out_{i}-$ 1 nodes, $|V|$ in the tree-lasso penalty is upper-bounded by $2out_{i}$.
Substituting (6) into (4) yields our sparse tree-guided lasso pruning with Information Bottleneck (sTLP-IB). Similarly, within each subtree formed by classes, we select the filters that could convey more information.
3.3 Optimization with Proximal Gradient Descent
With smoothing proximal gradient descent method Chen et al. (2012); Yuan (2024) we implemented, we could easily apply into (4) and yield the $\beta$ results for each layer, which also guarantee the convergence rate of O($\frac{1}{\epsilon}$) with polynomial time complexity, where $\epsilon$ is the differences between the yielded $\beta$ and the optimal $\beta$. We then prune the model $M$ based on the $\beta$ obtained from sGLP-IB and sTLP-IB.
@@ 3.3 Optimization with Proximal Gradient Descent
1:  for $i$ in $N$ do
2:     $\lambda_{lt}=\lambda_{l},\lambda_{ut}=\lambda_{u}$, $C_{i}\leftarrow M$.
@@
6:        $C_{i}^{*}$= sGLP-IB$(C_{i},\lambda)$ or sTLP-IB$(C_{i},\lambda)$.
7:        if  $size(C_{i}^{*})\leq E_{l}$  then
@@
9:        else if  $size(C_{i}^{*})\geq E_{u}$ then
@@
Table 1: Pruning results of VGG-16, ResNet-56, ResNet-110 and GoogleNet on CIFAR-10 dataset.
3.4 Adaptive Algorithm
We illustrate how our approach utilizes sGLP-IB and sTLP-IB to prune networks. We observe that they include a hyperparameter $\lambda$, which greatly influences the sparsity of $\beta$. If $\lambda$ is too small, each element of $\beta$ is a non-zero value, making the pruning ineffective; if $\lambda$ is too large, $\beta$ may contain too few non-zero values, making the pruning impossible as they will prune all the filters in one layer. Given that models have many layers, setting $\lambda$ for each layer can be very time-consuming. Therefore, we propose an adaptive method to selectively prune based on memory size and FLOPs requirements as shown in Algorithm 1.
In the adaptive method, we use a binary search approach. We set an upper and lower limit for the value of $\lambda$ and use exponential mapping to quickly find an appropriate $\lambda$ that fits the sparsity level of the structured Lasso. Given a range of the proportion of parameters to retain or a range of FLOPs to retain, our algorithm can adaptively find suitable $\lambda$ values for each layer to ensure that only a specific number of columns has non-zero values in $\beta$ and meet the requirements. This method ensures a relatively balanced number of parameters across layers, preventing any layer from having too many or too few parameters. It also allows for the setting of different parameter numbers and FLOP retention ratios for each layer.
4 Experiments
To demonstrate the effectiveness of our sGLP-IB and sTLP-IB, we conduct extensive experiments based on many representative CNNs, including various variants of VGGNet, ResNet, and GoogleNet, using three widely used datasets: CIFAR-10, CIFAR-100, and ImageNet. We also perform numerous ablation studies to analyze the robustness and characteristics of our methods.
[…]
@@ 4 Experiments
Baselines: Our extensive baselines are primarily based on the aforementioned structured pruning methods in Section 2.0.1. Besides, we include NORTON Pham et al. (2024a, b), which represents state-of-the-art unstructured pruning methods through the use of tensor decomposition.
4.1 Experiments on CIFAR-10
VGG-16. Table 1 presents our experiments on CIFAR-10 using VGGNet-16. Our methods, sGLP-IB and sTLP-IB, demonstrate strong performance, achieving Top-1 accuracy rates of 94.05% and 94.10%, and pruning ratios of 61% for FLOPs and 85% for parameters. sTLP-IB achieves the highest Top-1 accuracy among all previous baselines but with lower computational overhead and memory requirements, which is 0.14% higher than the pretrained model and demonstrates the effectiveness of utilizing class-wise information. Although sGLP-IB’s Top-1 accuracy is slightly lower than AutoBot by 0.05%, it achieves greater reductions in both FLOPs and parameters. We also present additional results based on varying pruning ratios, showing that our methods remain superior to most baselines even at higher pruning levels. This proves the effectiveness of using graph-structured lasso and tree-guided lasso, which utilize class-wise information, in pruning. Notably, sTLP-IB performs better than sGLP-IB, as the tree-structured approach, which constructs subtrees, captures class information more precisely than graphs.
ResNet-56. Table 1 also shows that our methods, sGLP-IB and sTLP-IB, outperform most baselines using ResNet-56 with Top-1 accuracy rates of 93.93% and 93.96%, respectively, while maintaining the same pruning ratios of 53% for FLOPs and 52% for parameters. sTLP-IB outperforms sGLP-IB, consistent with the results observed with VGGNet-16. Although NORTON slightly surpasses our methods by 0.04%, its pruning ratios were lower, with only 42% for FLOPs and 48% for parameters. Other methods like DECORE, FTWT, and FSW exhibit higher pruning ratios than SGLP-IB and STLP-IB, but they compromise Top-1 accuracy. Therefore, our methods achieve a balance between accuracy and pruning ratios, improving upon the pretrained model by 0.68%.
@@ 4.1 Experiments on CIFAR-10
ResNet-110. Both sGLPIB and sTLP-IB achieve state-of-the-art performance on ResNet-110. sTLP-IB behaves better with a Top-1 accuracy of 93.80%, surpassing the best baseline by 0.12%, while reducing 62% of FLOPs and 68% of parameters as Table 1, showing an improvement of 0.30% over the unpruned model. Notably, even though we choose the same pruning ratios, sGLP-IB eliminates more FLOPs and also surpasses all baselines with an accuracy of 93.69%, albeit 0.11% lower than sTLP-IB. This further underscores the effectiveness of the tree-guided lasso, which retains more critical filters, explaining the FLOPs difference compared to sGLP-IB. Our experiments on ResNet-56 and ResNet-110 validate the efficacy of our two pruning methods on ResNet architectures.
GoogleNet. On GoogleNet, our methods also demonstrate effectiveness, with sGLP-IB reaching 94.75% and sTLP-IB achieving a test accuracy of 94.78%. Despite removing 69% of FLOPs and 63% of parameters, we significantly reduce overhead and memory size. However, due to the Inception structure of GoogleNet, more fine-grained pruning is necessary. EPruner, which employs Affinity Propagation for fine-grained clustering, achieves better results but with a larger computational overhead compared to ours. Our methods surpass the remaining baselines in terms of accuracy, FLOPs pruning ratio, and parameter pruning ratio. The integration of class-wise information with our methods and other approaches holds substantial promise for future advancements.
4.2 Experiments on ImageNet
Compared to the CIFAR dataset, as shown in Table 2, we conduct experiments on a larger dataset, ImageNet, with numerous baselines using ResNet-50. The results indicate that our sTLP-IB achieves state-of-the-art performance, outperforming all baselines with a Top-1 accuracy of 76.12%, a 57% reduction in FLOPs, and a 55% reduction in parameters, with a mere 0.03% drop compared to the pretrained model. Our sGLP-IB also outperforms the majority of baselines and was comparable to CSD. Although sGLP-IB achieves a higher pruning ratio than CSD, its accuracy is only 0.01% lower, which may be attributed to CSD benefiting from its Discrete Wavelet Transform on large datasets, which shows the effectiveness of sGLP-IB. Combined with the results on CIFAR-10 from Section 4.1, our methods are applicable not only to simple datasets but also to complex, real-world datasets, showcasing their significant potential. We can leverage class-wise information and structured lasso for pruning across datasets.
4.3 Ablation Studies
Influence of batch sizes. As we keep the same pruning ratio as Table 1, the accuracy of sGLP-IB and sTLP-IB increases with the batch size but essentially converges once the batch size reaches 128 in Table 3. This observation forms the basis for setting our experimental batch size to 128. With a batch size of 64, the randomness due to smaller data volume results in sGLP-IB slightly outperforming sTLP-IB by 0.01% in Top-1 accuracy. As mentioned in work Lin et al. (2020), the average rank of feature maps generated by a single filter remains consistently the same and the pruned model’s accuracy is not significantly affected by the batch size during the pruning process. Nonetheless, sTLP-IB consistently outperforms sGLP-IB across varying batch sizes, aligning with our previous findings that tree-guided lasso is more effective than graph-structured lasso when the batch size is larger. The experimental results demonstrate that our methods are robust to batch size variations, provided that the random selection of samples within batches is maintained.
@@ 4.3 Ablation Studies
Figure 2: t-SNE visualizations of pruned models with different kernel functions before finetuning using VGGNet-16 on CIFAR-10 dataset, where T denotes sTLP-IB and G denotes sGLP-IB.
@@
t-SNE of the pruned model before finetuning. We further analyze the effectiveness of our method using t-SNE visualization on the pruned model before fine-tuning, following the previous settings. As illustrated in Figure 2, we also compare our approach with the most similar state-of-the-art method, APIB. Under the linear function, both sGLP-IB and sTLP-IB from our method perform better than APIB (Row 1), with more distinct clusters between classes, indicating that our approach better leverages class-wise information. We also present t-SNE visualizations for the Gaussian, Sigmoid, and Laplacian kernels (Row 2 and 3). We find that both the Gaussian and Laplacian kernels aid in learning class information, while the Sigmoid kernel function results in almost random distributions, consistent with the results in Table 5. We also evaluate the pruned model’s Top-1 accuracy before finetuning, as shown in Table 6. We observe that the Gaussian and Laplacian kernels facilitate learning more information in high-dimensional spaces, while the linear kernel demonstrates the inherent capability of our structured lasso. The results with the Sigmoid kernel mirror the t-SNE visualizations. Notably, both in Table 6 and Figure 2, sTLP-IB consistently outperformed sGLP-IB, aligning with the results in Table 1.
@@
Influence of pruning ratio. By adjusting the pruning ratio, we find that when the ratio is below 50%, the performance of sGLP-IB and sTLP-IB is similar due to the low pruning rate, resulting in nearly identical retained information and thus similar accuracy. When the ratio is below 70%, the variations in accuracy are minimal, indicating that our methods effectively capture and utilize statistical information for pruning. When the ratio falls below 10%, the accuracy has a slight drop, due to the excessive number of redundant filters.
@@
Influence of layer pruning. We also conduct experiments to assess the effect of retaining the initial few layers without pruning. To maintain the same pruning ratio, we compensate by pruning more in the later layers for a similar number of parameters. In this scenario, the accuracies for sGLP-IB and sTLP-IB are 93.76% and 93.88% on CIFAR-10 with VGGNet-16, respectively. If the first two layers were not pruned, the accuracies are 93.79% and 93.98%, respectively. These results are lower than those in Table 1, which can be attributed to the fact that the early layers focus on low-level features in the images, while the later layers contain more class-wise information.
Supplementary Experiments. Experiments on CIFAR-100 dataset, experiments with FLOPs pruning ratio and time complexity discussions can be found in Appendix.
5 Conclusion

### v2

@@ Introduction
Where each group of $j$-th subtree regression factors $\bm{\beta}_{j}^{G_{v}}$ is weighted by $w_{ve}$, as in (7), where $A$ is ancestor set. Vector $\bm{\beta}_{j}^{G_{v}}$ represents regression factors $\{\bm{\beta}_{j}^{k}\mid k\in G_{v}\}$. Each node $v\in V$ of one tree is associated with a group $G_{v}$, whose members are columns of $\bm{\beta}$ within nodes of the same subtree.
@@
Generally, $h_{v}$ represents the weight for selecting covariates relevant specifically to column of $\bm{\beta}$ linked to each child of node $v$, while $1-h_{v}$ is the weight for jointly selecting covariates relevant to all children of node $v$, $h_{v}\in(0,1)$. Assuming $|V|$ represents the number of nodes in a tree, since a tree associated with $out_{i}$ columns of $\bm{\beta}$ for $\bm{C}_{i}$ has $2out_{i}-1$ nodes, $|V|$ in tree-lasso penalty is upper-bounded by $2out_{i}$.
Substituting (6) into (4) yields our sparse tree-guided lasso pruning with Information Bottleneck (sTLP-IB) method. Similarly, within each subtree formed by classes, we select the filters that could convey more information.
Optimization with Proximal Gradient Descent
With smoothing proximal gradient descent method (Chen et al. 2012; Yuan 2024) we implemented, we could easily apply into (4) and yield the $\bm{\beta}$ results for each layer, which also guarantee the convergence rate of O($\frac{1}{\epsilon}$) with polynomial time complexity, where $\epsilon$ is the differences between the yielded $\bm{\beta}$ and the optimal $\bm{\beta}^{*}$. We then prune the model $M$ based on the $\bm{\beta}$ obtained from sGLP-IB and sTLP-IB.
@@
1:for $i$ in NN do
2:  $\lambda_{lt}=\lambda_{l},\lambda_{ut}=\lambda_{u}$, $\bm{C}_{i}\leftarrow M$.
@@
[…]
7:   if $size(\bm{C}_{i}^{*})\leq E_{l}$ then
@@
9:   else if $size(\bm{C}_{i}^{*})\geq E_{u}$ then
@@
Adaptive Algorithm
We illustrate how our approach utilizes sGLP-IB and sTLP-IB to prune networks. We observe that they include a hyperparameter $\lambda$, which greatly influences the sparsity of $\bm{\beta}$. If $\lambda$ is too small, each element of $\bm{\beta}$ is a non-zero value, making the pruning ineffective; if $\lambda$ is too large, $\bm{\beta}$ may contain too few non-zero values, making the pruning impossible as they will prune all the filters in one layer. Given that models have many layers, setting $\lambda$ for each layer can be very time-consuming. Therefore, we propose an adaptive method to selectively prune based on memory size and FLOPs requirements as shown in Algorithm 1.
In the adaptive method, we use a binary search approach. We set an upper and lower limit for the value of $\lambda$ and use exponential mapping to quickly find an appropriate $\lambda$ that fits the sparsity level of the structured Lasso. Given a range of the proportion of parameters to retain or a range of FLOPs to retain, our algorithm can adaptively find suitable $\lambda$ values for each layer to ensure that only a specific number of columns have non-zero values in $\bm{\beta}$ and meet the requirements. This method ensures a relatively balanced number of parameters across layers, preventing any layer from having too many or too few parameters. It also allows for the setting of different parameter numbers and FLOP retention ratios for each layer.
@@
Table 1: Pruning results of VGG-16, ResNet-56, ResNet-110 and GoogleNet on CIFAR-10 dataset. For VGG-16, we present results under different pruning ratios.
Experiments
To demonstrate the effectiveness of our sGLP-IB and sTLP-IB, we conduct extensive experiments based on many representative CNNs, including various variants of VGGNet, ResNet, and GoogleNet, using three widely used datasets: CIFAR-10, CIFAR-100, and ImageNet. We also conduct experiments on resource-constrained devices (Raspberry Pi 4) and perform numerous ablation studies to analyze the robustness and characteristics of our methods.
Implementation Details: All experiments are implemented using PyTorch and averaged over five trials with random seeds. We use the SGD optimizer with a decaying learning rate initialized to 1ee-4, and we set the batch size to 128. For sGLP-IB, we choose a threshold of 0.61,8 as graph-structured lasso. After pruning, for fair comparison, we finetune VGGNet, ResNet and GoogleNet for 400 to 500 epochs. On ImageNet datasets, we train VGGNet and ResNet for 150 to 200 epochs. We choose the Laplacian kernel as our base kernel function for Gram matrix. We prune without the first four layers.
@@
Baselines: Our extensive baselines are primarily based on the aforementioned structured pruning methods in Section Structured Pruning. Besides, to provide a more comprehensive evaluation, we include NORTON (Pham, Zniyed, and Nguyen 2024a, b), which represents state-of-the-art unstructured pruning methods through the use of tensor decomposition.
Experiments on CIFAR-10
VGG-16. Table 1 presents our experiments on CIFAR-10 using VGGNet-16. Our methods, sGLP-IB and sTLP-IB, demonstrate strong performance, achieving accuracy of 94.05% and 94.10%, and pruning ratios of 61% for FLOPs and 85% for parameters. sTLP-IB achieves the highest Top-1 accuracy among all previous baselines with lower computational overhead and memory requirements, which is 0.14% higher than the pretrained model and demonstrates the effectiveness of utilizing structured class-wise information. Although sGLP-IB’s Top-1 accuracy is slightly lower than AutoBot by 0.05%, it achieves greater reductions in both FLOPs and parameters. We also present additional results based on varying pruning ratios, showing that our methods remain superior to most baselines even at higher pruning levels. This proves the effectiveness of using graph-structured lasso and tree-guided lasso, which utilize class-wise information, in pruning. Notably, sTLP-IB performs better than sGLP-IB, as the tree-structured approach, which constructs subtrees, captures class information more precisely than graphs.
ResNet-56. Table 1 also shows our methods, sGLP-IB and sTLP-IB, outperform most baselines using ResNet-56 with Top-1 accuracy rates of 93.93% and 93.96%, respectively, while maintaining same pruning ratios of 53% for FLOPs and 52% for parameters. sTLP-IB outperforms sGLP-IB, consistent with the results observed with VGGNet-16. Although NORTON slightly surpasses our methods by 0.04%, its pruning ratios were lower, with only 42% for FLOPs and 48% for parameters. Others like DECORE, FTWT, and FSW exhibit higher pruning ratios than SGLP-IB and STLP-IB, but they compromise accuracy. Therefore, our methods achieve a balance between accuracy and pruning ratios, improving upon the pretrained model by 0.68%.
@@
ResNet-110. Both sGLPIB and sTLP-IB achieve state-of-the-art performance on ResNet-110. sTLP-IB behaves better with accuracy of 93.80%, surpassing the best baseline by 0.12%, while reducing 62% of FLOPs and 68% of parameters as Table 1, showing an improvement of 0.30% over the unpruned model. Notably, even though we choose the same pruning ratios, sGLP-IB eliminates more FLOPs and also surpasses all baselines with an accuracy of 93.69%, albeit 0.11% lower than sTLP-IB. This further underscores the effectiveness of tree-guided lasso, which retains more critical filters, explaining the FLOPs difference compared to sGLP-IB. Our experiments on ResNet-56 and ResNet-110 validate the efficacy of our pruning methods on ResNet architectures.
GoogleNet. On GoogleNet, our methods also demonstrate effectiveness, with sGLP-IB reaching 94.75% and sTLP-IB achieving a test accuracy of 94.78%. Despite removing 69% of FLOPs and 63% of parameters, we significantly reduce overhead and memory size. However, due to the Inception structure of GoogleNet, more fine-grained pruning is necessary. EPruner, which employs Affinity Propagation for fine-grained clustering, achieves better results but with a larger computational overhead compared to ours. Our methods surpass the remaining baselines in terms of accuracy, FLOPs pruning ratio, and parameter pruning ratio. The integration of structured class-wise information with our methods and other approaches holds substantial promise for future advancements.
Experiments on ImageNet
Compared to the CIFAR dataset, as shown in Table 2, we conduct experiments on a larger dataset, ImageNet, with numerous baselines using ResNet-50. The results indicate that our sTLP-IB achieves state-of-the-art performance, outperforming all baselines with a Top-1 accuracy of 76.12%, a 57% reduction in FLOPs, and a 55% reduction in parameters, with a mere 0.03% drop compared to the pretrained model. Our sGLP-IB also outperforms the majority of baselines and was comparable to CSD. Although sGLP-IB achieves a higher pruning ratio than CSD, its accuracy is only 0.01% lower, which may be attributed to CSD benefiting from its Discrete Wavelet Transform on large datasets, which again shows the effectiveness of our sGLP-IB. Combined with the results on CIFAR-10 from Section Experiments on CIFAR-10, our methods are applicable not only to simple datasets but also to complex, real-world datasets, showcasing their significant potential. We can leverage class-wise information and structured lasso for pruning across datasets
@@
Ablation Studies
Influence of batch sizes. As we keep the same pruning ratio as Table 1, the accuracy of sGLP-IB and sTLP-IB increases with the batch size but essentially converges once the batch size reaches 128 in Table 3. This observation forms the basis for setting our experimental batch size to 128, and also ensures a fair comparison with other methods. With a batch size of 64, the randomness due to smaller data volume results in sGLP-IB slightly outperforming sTLP-IB by 0.01% in Top-1 accuracy. As mentioned in work (Lin et al. 2020), the average rank of feature maps generated by a single filter remains consistently the same and the pruned model’s accuracy is not significantly affected by the batch size during the pruning process. Nonetheless, sTLP-IB consistently outperforms sGLP-IB across varying batch sizes, aligning with our previous findings that tree-guided lasso is more effective than graph-structured lasso when the batch size is larger. The experimental results demonstrate that our methods are robust to batch size variations, provided that the random selection of samples within batches is maintained.
@@
Figure 2: t-SNE visualizations of pruned models with different kernel functions before finetuning using VGGNet-16 on CIFAR-10 (T denotes sTLP-IB and G denotes sGLP-IB).
@@
t-SNE of the pruned model before finetuning. We further analyze the effectiveness of our method using t-SNE visualization on the pruned model before fine-tuning, following the previous settings. As illustrated in Fig. 2, we also compare our approach with the most similar state-of-the-art method, APIB. Under the linear function, both sGLP-IB and sTLP-IB from our method perform better than APIB (Row 1), with more distinct clusters between classes, indicating that our approach better leverages structured class-wise information. We also present t-SNE visualizations for the Gaussian, Sigmoid, and Laplacian kernels (Row 2 and Row 3). We find that both the Gaussian and Laplacian kernels aid in learning class information, while the Sigmoid kernel function results in almost random distributions, consistent with the results in Table 5. We also evaluate the pruned model’s Top-1 accuracy before finetuning, as shown in Table 6. We observe that the Gaussian and Laplacian kernels facilitate learning more information in high-dimensional spaces, while the linear kernel demonstrates the inherent capability of our structured lasso. The results with the Sigmoid kernel mirror the t-SNE visualizations. Notably, both in Table 6 and Fig. 2, sTLP-IB consistently outperformed sGLP-IB, which aligns with the results in Table 1.
@@
Influence of pruning ratio for Parameters. By adjusting the pruning ratio in Table 7, we find that when the ratio is below 50%, the performance of sGLP-IB and sTLP-IB is similar due to the low pruning rate, resulting in nearly identical retained information and thus similar accuracy. When the ratio is below 70%, the variations in accuracy are minimal, indicating that our methods effectively capture and utilize statistical information for pruning. When the ratio falls below 10%, the accuracy has a slight drop, due to the excessive number of redundant filters.
@@
Influence of layer pruning. We conduct experiments to assess effects of retaining the initial few layers without pruning. To maintain the same pruning ratio, we compensate by pruning more in later layers. In this scenario, accuracies for sGLP-IB and sTLP-IB are 93.76% and 93.88% on CIFAR-10 with VGGNet-16, respectively. If the first two layers were not pruned, the accuracies are 93.79% and 93.98%, respectively. These results are lower than those in Table 1, which can be attributed to the fact that early layers focus on low-level features of inputs, while later layers contain more structured class-wise information.
Supplementary Materials
Artifact evaluation details, time complexity and convergence discussions, adaption for pruning ResNet and GoogleNet, dataset details, experiments with CIFAR-100 dataset, experiments with pruning ratios for FLOPs, real implementation on resource-constrained devices and our resource-constrained device prototype can be found in Appendix.
Discussions
Our paper demonstrates the effectiveness of utilizing structured class-wise information. Although this work focuses on classification, our method can also be applied to segmentation and object detection in the future, as filters likewise exhibit structured class information. Therefore, exploring better methods to capture more precise class-wise information is one of our future directions. Our method is orthogonal to quantization approaches (Li et al. 2023; Li and Gu 2023), and can be combined with them in the future to further reduce computational overhead. Moreover, as Transformer (Vaswani et al. 2017) architectures are increasingly adopted in multimedia applications and their large model sizes pose challenges for deployment on resource-constrained devices, we plan to explore applying our effective method to Transformer-based models as our future research focus to explore the effectiveness of our methods.
Conclusion